OpenStudy (anonymous):

12cm/d= how many cm per hour? Medals for simplest answer!

OpenStudy (anonymous):

(12cm/d)(1d/24h)=(12/24)cm/h=(1/2)cm/h Dimensional analysis

OpenStudy (anonymous):

I'm sorry I don't understand that